# Hessian stability and convergence rates for entropic and Sinkhorn potentials via semiconcavity

### v1

@@ 2.1. Entropic and gradients’ stability
This directly follows from [45, Theorem 2.1.b] after noticing that
$\frac{\mathrm{d}\pi^{\mu}}{\mathrm{d}\pi^{\nu}}=\exp((\varphi^{\nu}-\varphi^{\mu})\oplus(\psi^{\nu}-\psi^{\mu}))\quad\mathrm{R}_{0T}-a.s.$
and hence also $\pi^{\nu}$-a.s. (since $\mathscr{H}(\pi^{\nu}|\mathrm{R}_{0T})<\infty$). Notice that [45, Theorem 2.1.b] further implies $\mathscr{H}(\pi^{\mu}|\pi^{\nu})<\infty$.
We now proceed to bound $\mathscr{H}(\pi^{\mu}|\pi^{\nu})$ exhibiting a suitable admissible plan in (2.9). In view of that, let us consider a transport map $\mathcal{T}$ from $\mu$ to $\nu$, that is such that $\mathcal{T}_{\#}\mu=\nu$ and take $X_{0}\sim\mu$ and define $X_{0}^{\nu}\coloneqq\mathcal{T}(X_{0})\sim\nu$. The existence of such map follows from [51, Corollary 1.29] since $\mu\ll\mathrm{Leb}$. For any $x$ we also define the process $(X^{\varphi^{\nu},x}_{s})_{s\in[0,T]}$ by
(2.10) $\mathrm{d}X^{\varphi^{\nu},x}_{s}=-\nabla\varphi^{\nu}_{s}(X^{\varphi^{\nu},x}_{s})\mathrm{d}s+\mathrm{d}B_{s}\,,\quad X^{\varphi^{\nu},x}_{0}=x,$
where $B_{\cdot}$ is a Brownian motion independent of $(X_{0},X^{\nu}_{0})$. Finally, let $\gamma_{s}=(1-s/T)X_{0}+(s/T)X^{\nu}_{0}$ and consider now the stochastic process $X_{\cdot}$ given by
$X_{s}=X^{\varphi^{\nu},\gamma_{s}}_{s},\quad s\in[0,T]$
and note that if we call $\pi_{\mathrm{comp}}$ the law of $(X_{T},X_{0})$, then $\pi_{\mathrm{comp}}\in\Pi(\rho,\,\mu)$.
@@
Next, notice that $X_{T}=X_{T}^{\varphi^{\nu},X_{0}^{\nu}}=X_{T}^{\varphi^{\nu},\mathcal{T}(X_{0})}$ while $X_{0}=X_{0}^{\varphi^{\nu},X_{0}}=X_{0}$, and hence the conditional probabilities appearing in the last display are one the translation of the other, that is
@@
In particular, this combined with Section 2.1 implies for $\mu$-a.e. $z\in\mathbb{R}^{d}$ that
$\mathscr{H}(\pi_{\mathrm{comp}}(\cdot|z)|\pi^{\nu}(\cdot|z))=$ $\,\mathscr{H}(\pi^{\nu}(\cdot|\mathcal{T}(z))|\pi^{\nu}(\cdot|z))\leq\frac{\Lambda(\varphi^{\nu}_{0})}{2T}\,|\mathcal{T}(z)-z|^{2}\,,$
which integrated yields to
$\mathscr{H}(\pi^{\mu}|\pi^{\nu})\leq\mathscr{H}(\mu|\nu)+\frac{\Lambda(\varphi^{\nu}_{0})}{2T}\,\int|\mathcal{T}(z)-z|^{2}\,\mathrm{d}\mu(z).$
By minimizing the last display over all the measurable transport maps from $\mu$ to $\nu$ we conclude the proof of the entropic stability bound (see [51, Theorem 1.22] for the existence of the optimal transport map for $\mathbf{W}_{2}(\mu,\nu)$).
Let us now focus on the proof of (2.6). If in H 2 we assume $\mathscr{H}(\mu|\nu)<+\infty$, then the conclusion follows from the disintegration property of the relative entropy (see for instance [45, Lemma 1.6] and [40, Appendix A]) since
$\mathbb{E}_{\mu}[\mathscr{H}(\pi^{\mu}(\cdot|X)|\pi^{\nu}(\cdot|X))]=\mathscr{H}(\pi^{\mu}|\pi^{\nu})-\mathscr{H}(\mu|\nu)\,,$
which combined with the above entropic stability bound concludes the proof of (2.6) under a finite entropy assumption.
On the other hand, if we consider in H 2 the case $\mu\ll\nu$ with $\Lambda(\varphi^{\mu}_{0})$ finite (e.g. $\rho$ with compact support or log-concave density), then we can use an approximation argument and consider the sequence of probability measures $\mu^{n}\in\mathcal{P}(\mathbb{R}^{d})$ whose densities are defined as
@@
[…]
@@
By combining this last bound with Fatou’s lemma and with the entropic stability estimate already proven above we get
@@
[…]
@@
Finally, the proof of (2.7) follows from (2.6) since from Itô’s formula and (2.1) we immediately see that
@@
which combined with Grönwall’s lemma gives for any $s\leq\delta T$
$\mathbb{E}[|Y^{\eta}_{0}|^{2}]\,e^{\int_{0}^{s}2\lambda(\varphi^{\nu}_{t})\mathrm{d}t}\leq\mathbb{E}[|Y^{\eta}_{s}|^{2}]\,.$
When integrated over $s\in[0,T]$, this inequality reads as
$\mathbb{E}[|Y^{\eta}_{0}|^{2}]\leq\biggl{(}\int_{0}^{T}e^{\int_{0}^{s}2\lambda(\varphi^{\nu}_{t})\mathrm{d}t}\mathrm{d}s\biggr{)}^{-1}\int_{0}^{T}\mathbb{E}[|Y^{\eta}_{s}|^{2}]\mathrm{d}s=\frac{C^{\varphi^{\nu}}}{T}\int_{0}^{T}\mathbb{E}[|Y^{\eta}_{s}|^{2}]\mathrm{d}s\,,$
and combining it with the energy-entropic identity (2.3) gives
$\mathbb{E}[|Y^{\eta}_{0}|^{2}]\leq\frac{2C^{\varphi^{\nu}}}{T}\,\mathbb{E}_{\mu}[\mathscr{H}(\pi^{\mu}(\cdot|X)|\pi^{\nu}(\cdot|X))]\,.$
@@
Firstly, fix a coupling $\tau\in\Pi(\mu,\nu)$ between our two target marginals and consider the probability measures on $(\mathbb{R}^{d})^{3}$ defined by the densities $\pi^{\mu}_{s}(\mathrm{d}x|y)\tau(\mathrm{d}y,\mathrm{d}z)$ and $\pi^{\nu}_{s}(\mathrm{d}x|z)\tau(\mathrm{d}y,\mathrm{d}z)$ (for notations’ sake we indicate these two probabilities respectively with $\pi^{\mu}_{s}(\cdot|y)\otimes\tau$ and $\pi^{\nu}_{s}(\cdot|z)\otimes\tau$). Clearly we have
@@
therefore, from the data processing inequality and from the disintegration property of relative entropy (cf. [45, Lemma 1.6] and [40, Appendix A]) we deduce that
@@
$\mathscr{H}(\pi^{\mu}_{s}|\pi^{\nu}_{s})\leq\biggl{(}\frac{\Lambda(\varphi^{\nu}_{0})}{2T}+\frac{s^{-1}-T^{-1}}{2}\biggr{)}\int|z-y|^{2}\mathrm{d}\tau+\int\mathbb{E}[\eta_{s}(X^{\varphi_{\mu},y}_{s})-\eta_{0}(y)]\mathrm{d}\mu(y)$
$-\int\langle\nabla\eta_{0}(y),z-y\rangle\mathrm{d}\tau$
$\leq\biggl{(}\frac{\Lambda(\varphi^{\nu}_{0})}{2T}+\frac{s^{-1}-T^{-1}}{2}\biggr{)}\int|z-y|^{2}\mathrm{d}\tau+\mathbb{E}[\eta_{s}(X^{\varphi_{\mu},\mu}_{s})-\eta_{0}(X_{0}^{\varphi^{\mu},\mu})]$
$+\|\nabla\eta_{0}\|_{\mathrm{L}^{2}(\mu)}\biggl{(}\int|z-y|^{2}\mathrm{d}\tau\biggr{)}^{\nicefrac{{1}}{{2}}}\,.$
Now, since $\|\nabla\eta_{0}\|_{\mathrm{L}^{2}(\mu)}^{2}=\mathbb{E}[|Y^{\eta}_{0}|^{2}]$, from (2.7) we deduce that
$\mathscr{H}(\pi^{\mu}_{s}|\pi^{\nu}_{s})\leq\biggl{(}\frac{\Lambda(\varphi^{\nu}_{0})}{2T}+\frac{s^{-1}-T^{-1}}{2}\biggr{)}\int|z-y|^{2}\mathrm{d}\tau+\mathbb{E}[\eta_{s}(X^{\varphi_{\mu},\mu}_{s})-\eta_{0}(X_{0}^{\varphi^{\mu},\mu})]$
$+\frac{\sqrt{\Lambda(\varphi^{\nu}_{0})\,C^{\varphi^{\nu}}}}{T}\,\mathbf{W}_{2}(\mu,\nu)\,\biggl{(}\int|z-y|^{2}\mathrm{d}\tau\biggr{)}^{\nicefrac{{1}}{{2}}}\,.$
and by minimizing over $\tau\in\Pi(\mu,\nu)$, we obtain
@@
and hence that
[…]
which combined with Theorem 2.1 leads to our thesis. ∎
Notice that the above bound diverges as $s\downarrow 0$. This should not be surprising since for $s=0$ we are trying to bound $\mathscr{H}(\mu|\nu)$ solely with the Wasserstein distance (and the latter might still be finite while the former diverges). Therefore, the previous lemma shows that the contribution of $\mathscr{H}(\mu|\nu)$ is involved only for $s=0$, while as soon as $s>0$ the Wasserstein distance is enough.
We conclude this section with a useful integral bound which will be employed in the proof of our main result in the next section.
[…]
@@
In view of Girsanov’s Theorem identity (2.5), we clearly have
$\int_{0}^{\delta T}\mathbb{E}[|Y^{\theta}_{s}|^{2}]\mathrm{d}s=2\mathscr{H}(\mathcal{L}(X^{\psi^{\mu},\rho}_{[0,\delta T]})|\mathcal{L}(X^{\psi^{\nu},\rho}_{[0,\delta T]}))=\int_{(1-\delta)T}^{T}\mathbb{E}[|Y^{\eta}_{s}|^{2}]\mathrm{d}s+2\mathscr{H}(\mathcal{L}(X^{\varphi^{\mu},\mu}_{(1-\delta)T})|\mathcal{L}(X^{\varphi^{\nu},\nu}_{(1-\delta)T}))$
$\leq\int_{0}^{T}\mathbb{E}[|Y^{\eta}_{s}|^{2}]\mathrm{d}s+2\mathscr{H}(\mathcal{L}(X^{\varphi^{\mu},\mu}_{(1-\delta)T})|\mathcal{L}(X^{\varphi^{\nu},\nu}_{(1-\delta)T}))$
=(2.3)${=}2\mathbb{E}_{\mu}[\mathscr{H}(\pi^{\mu}(\cdot|X)|\pi^{\nu}(\cdot|X))]+2\mathscr{H}(\mathcal{L}(X^{\varphi^{\mu},\mu}_{(1-\delta)T})|\mathcal{L}(X^{\varphi^{\nu},\nu}_{(1-\delta)T}))$ )
where we have relied on a second application of Girsanov’s Theorem (as we did for (2.4)), combined with the time-reversal identities (1.5). By applying Theorem 2.1 and Section 2.1 we finally conclude that
3. Proof of main results
@@ 3.1. Quantitative stability estimates of gradients
Assume H 1 and H 2, fix $\delta\in(0,1)$, and let
$C^{\psi^{\nu}}_{\delta}\coloneqq T\biggl{(}\int_{0}^{\delta T}e^{\int_{0}^{s}2\lambda(\psi^{\nu}_{t})\mathrm{d}t}\mathrm{d}s\biggr{)}^{-1}\,.$
Then we have
@@
As a corollary, if we define
$C_{\rho\nu}^{\delta}\coloneqq C^{\psi^{\nu}}_{\delta}\,\biggl{(}\frac{\delta}{1-\delta}+3\,\Lambda(\varphi^{\nu}_{0})+2\sqrt{\Lambda(\varphi^{\nu}_{0})\,C^{\varphi^{\nu}}}\biggr{)}\,,$
then we have
@@
From Itô’s formula and (2.2), for all $s\leq\delta T$ we have
@@
which combined with Grönwall’s lemma gives for any $s\leq\delta T$
$\mathbb{E}[|Y^{\theta}_{0}|^{2}]\,e^{\int_{0}^{s}2\lambda(\psi^{\nu}_{t})\mathrm{d}t}\leq\mathbb{E}[|Y^{\theta}_{s}|^{2}]\,,$
that integrated over $s\in[0,\delta T]$ reads as
@@
Next, notice that from Girsanov’s theory (namely, the energy entropy identity (2.5)) we may recognize in the above right-hand side the relative entropy on the path space between the Schrödinger bridge from $\rho$ to $\mu$ and the Schrödinger bridge from $\rho$ to $\nu$, restricted on the time interval $[0,\delta T]$, that is
@@
By recalling the time-reversal identities (1.5) and by applying the disintegration property of relative entropies (cf. [45, Lemma 1.6] and [40, Appendix A]) and Girsanov’s Theorem (w.r.t. the backward corrector process $Y^{\eta}_{\cdot}$) we deduce that
@@
This proves our first claim.
By recalling the identity (2.3) we then have
@@
which can be bounded with Section 2.1 and with Theorem 2.1, yielding to
$\mathbb{E}[|Y^{\theta}_{0}|^{2}]\leq\frac{C^{\psi^{\nu}}_{\delta}}{T}\biggl{(}\frac{2\,\Lambda(\varphi^{\nu}_{0})}{T}+T^{-1}\biggl{(}\frac{1}{1-\delta}-1\biggr{)}+2\frac{\sqrt{\Lambda(\varphi^{\nu}_{0})\,C^{\varphi^{\nu}}}}{T}\biggr{)}\mathbf{W}_{2}^{2}(\mu,\nu)\,.$
Hereafter we specify Theorem 3.1 to two different settings. What follows is based on the explicit computations performed in Appendix A and Appendix B. From a notational viewpoint, we will write $a\lesssim b$ whenever there exists a numerical constant $C>0$ (independent of $T,\nu,\rho,\mu$) such that $a\leq C\,b$.
[…]
Assume H 1, that $\mu\ll\nu$, and that both $\rho$ and $\nu$ are compactly supported in a ball of radius $R$ (big enough so that $R^{2}\geq T$). Then we have
Proof.
Since $\rho$ has compact support, the same computations performed in Section A.1 guarantee $\Lambda(\varphi^{\mu}_{0})<\infty$ and hence the validity of H 2. Our choice of $\delta$ in (B.2) and the following computations yield to $C_{\rho\nu}^{\delta}\lesssim R^{4}\,T^{-2}$. ∎
[…]
Assume H 1, that $\mu\ll\nu$, and that both $\rho$ and $\nu$ are log-concave, i.e., that their (negative) log-densities satisfy $\nabla^{2}U_{\rho}\geq\alpha_{\rho}$ and $\nabla^{2}U_{\nu}\geq\alpha_{\nu}$ for some $\alpha_{\rho},\,\alpha_{\nu}>0$ (w.l.o.g. such that $\alpha_{\rho}\vee\alpha_{\nu}<T^{-1}$). Then we have
Proof.
Since $\rho$ has log-concave density, the same computations performed in Section A.2 guarantee $\Lambda(\varphi^{\mu}_{0})<\infty$ and hence the validity of H 2. Our choice of $\delta$ in (B.2) and the following computations yield to $C_{\rho\nu}^{\delta}\lesssim\alpha_{\rho}^{-1}\,\alpha_{\nu}^{-1}\,T^{-2}$. ∎
As already mentioned in the introduction, our explicit computations allow to straightforwardly deduce bounds akin Section 3.1 and Section 3.1 when considering one marginal log-concave and the other one with compact support.

### v2

@@ 2.1. Entropic and gradients’ stability
This directly follows from [Nut21, Theorem 2.1.b] after noticing that almost surely $\frac{\mathrm{d}\pi^{\mu}}{\mathrm{d}\pi^{\nu}}=\exp((\varphi^{\nu}-\varphi^{\mu})\oplus(\psi^{\nu}-\psi^{\mu}))$ and hence also $\pi^{\nu}$-a.s.  Notice that [Nut21, Theorem 2.1.b] further implies $\mathscr{H}(\pi^{\mu}|\pi^{\nu})<\infty$. We now proceed to bound $\mathscr{H}(\pi^{\mu}|\pi^{\nu})$ exhibiting a suitable admissible plan in (2.9). In view of that, let us consider the optimal transport map $\mathcal{T}$ from $\mu$ to $\nu$, that is such that $\mathcal{T}_{\#}\mu=\nu$ (see [San15, Theorem 1.22] for the existence of the optimal transport map for $\mathbf{W}_{2}(\mu,\nu)$). Next, for any $x\in\mathbb{R}^{d}$ consider the backward process $(X^{\varphi^{\nu},x}_{s})_{s\in[0,T]}$ defined by (2.8), and independently take $X_{0}\sim\mu$ and define $X_{0}^{\nu}\coloneqq\mathcal{T}(X_{0})\sim\nu$. Finally, let $\gamma_{s}=(1-s/T)X_{0}+(s/T)X^{\nu}_{0}$ and consider now the stochastic process $X_{\cdot}$ defined for any $s\in[0,T]$ as $X_{s}=X^{\varphi^{\nu},\gamma_{s}}_{s}$, and note that if we call $\pi_{\mathrm{comp}}$ the law of $(X_{T},X_{0})$, then $\pi_{\mathrm{comp}}\in\Pi(\rho,\,\mu)$.
@@
Next, notice that $X_{T}=X_{T}^{\varphi^{\nu},X_{0}^{\nu}}=X_{T}^{\varphi^{\nu},\mathcal{T}(X_{0})}$ while $X_{0}=X_{0}^{\varphi^{\nu},X_{0}}=X_{0}$, and hence the conditional probabilities appearing in the last display are translations, that is
@@
This combined with (2.10) and Section 2.1 proves our claim since the latter implies
$\int\mathscr{H}(\pi^{\nu}(\cdot|\mathcal{T}(z))|\pi^{\nu}(\cdot|z))\mathrm{d}\mu(z)\leq\frac{\Lambda(\varphi^{\nu}_{0})}{2T}\int|\mathcal{T}(z)-z|^{2}\mathrm{d}\mu(z)=\frac{\Lambda(\varphi^{\nu}_{0})}{2T}\,\mathbf{W}_{2}^{2}(\mu,\nu)\,.$
Let us now focus on the proof of (2.6a). If in H 2 we assume $\mathscr{H}(\mu|\nu)<+\infty$, then the conclusion follows from the disintegration property of the relative entropy (see for instance [Nut21, Lemma 1.6] and [Léo14, Appendix A]) since $\mathbb{E}_{\mu}[\mathscr{H}(\pi^{\mu}(\cdot|X)|\pi^{\nu}(\cdot|X))]=\mathscr{H}(\pi^{\mu}|\pi^{\nu})-\mathscr{H}(\mu|\nu)$, which combined with the above entropic stability bound concludes the proof of (2.6a) under a finite entropy assumption.
On the other hand, if we assume that $\mu\ll\nu$ with $\Lambda(\varphi^{\mu}_{0})$ finite (e.g. $\rho$ with compact support or log-concave density), then we can use an approximation argument and consider the sequence of probability measures $\mu^{n}\in\mathcal{P}(\mathbb{R}^{d})$ whose densities are defined as
@@
$\mathscr{H}(\mu^{n}|\mu)=-\log(C_{n})+\int\log\biggl(\mathbf{1}_{\{\nicefrac{{\mathrm{d}\mu}}{{\mathrm{d}\nu}}\leq n\}}+n\mathbf{1}_{\{\nicefrac{{\mathrm{d}\mu}}{{\mathrm{d}\nu}}>n\}}\frac{\mathrm{d}\nu}{\mathrm{d}\mu}\biggr)\mathrm{d}\mu^{n}$
$\leq-\log(C_{n})+\int n\mathbf{1}_{\{\nicefrac{{\mathrm{d}\mu}}{{\mathrm{d}\nu}}>n\}}\frac{\mathrm{d}\nu}{\mathrm{d}\mu}\,\mathrm{d}\mu^{n}\leq 1-\log(C_{n})<+\infty\,.$
@@
By combining this last bound with Fatou’s lemma and with the entropic stability estimate already proven above we finally get
@@
$=\liminf_{n\to\infty}\,\mathbb{E}_{\mu^{n}}[\mathscr{H}(\pi^{\mu^{n}}(\cdot|X)|\pi^{\nu}(\cdot|X))]=\liminf_{n\to\infty}\,\mathscr{H}(\pi^{\mu^{n}}|\pi^{\nu})-\mathscr{H}(\mu^{n}|\nu)$
@@
Finally, the proof of (2.6b) follows from (2.6a) since from Itô’s formula and (2.1a) we see that
@@
which combined with Grönwall’s lemma gives $\mathbb{E}[|Y^{\eta}_{0}|^{2}]\,e^{\int_{0}^{s}2\lambda(\varphi^{\nu}_{t})\mathrm{d}t}\leq\mathbb{E}[|Y^{\eta}_{s}|^{2}]$. When integrated over $s\in[0,T]$, this inequality reads as
$\mathbb{E}[|Y^{\eta}_{0}|^{2}]\leq\biggl(\int_{0}^{T}e^{\int_{0}^{s}2\lambda(\varphi^{\nu}_{t})\mathrm{d}t}\mathrm{d}s\biggr)^{-1}\int_{0}^{T}\mathbb{E}[|Y^{\eta}_{s}|^{2}]\mathrm{d}s=\frac{C^{\varphi^{\nu}}}{T}\int_{0}^{T}\mathbb{E}[|Y^{\eta}_{s}|^{2}]\mathrm{d}s\,.$
Our thesis then follows by combining this last bound with (2.2) and (2.6a). ∎
@@
Firstly, let $\tau\in\Pi(\mu,\nu)$ be the optimal transport coupling between our two target marginals and consider the probability measures on $(\mathbb{R}^{d})^{3}$ defined by the densities $\pi^{\mu}_{s}(\mathrm{d}x|y)\tau(\mathrm{d}y,\mathrm{d}z)$ and $\pi^{\nu}_{s}(\mathrm{d}x|z)\tau(\mathrm{d}y,\mathrm{d}z)$ (for notations’ sake we indicate these two probabilities respectively with $\pi^{\mu}_{s}(\cdot|y)\otimes\tau$ and $\pi^{\nu}_{s}(\cdot|z)\otimes\tau$).
Clearly we have
@@
therefore, from the data processing inequality and from the disintegration property of relative entropy (cf. [Nut21, Lemma 1.6] and [Léo14, Appendix A]) we deduce that
@@
$\mathscr{H}(\pi^{\mu}_{s}|\pi^{\nu}_{s})\leq\biggl(\frac{\Lambda(\varphi^{\nu}_{0})}{2T}+\frac{s^{-1}-T^{-1}}{2}\biggr)\mathbf{W}_{2}^{2}(\mu,\nu)+\mathbb{E}[\eta_{s}(X^{\varphi_{\mu},\mu}_{s})-\eta_{0}(X_{0}^{\varphi^{\mu},\mu})]+\|\nabla\eta_{0}\|_{\mathrm{L}^{2}(\mu)}\,\mathbf{W}_{2}(\mu,\nu)\,.$
Now, since $\|\nabla\eta_{0}\|_{\mathrm{L}^{2}(\mu)}^{2}=\mathbb{E}[|Y^{\eta}_{0}|^{2}]$, from (2.6b) we deduce that
@@
which combined with (2.2) and (2.6a) finally gives
$\mathbb{E}[\eta_{s}(X^{\varphi_{\mu},\mu}_{s})-\eta_{0}(X^{\varphi^{\mu},\mu}_{0})]=\frac{1}{2}\int_{0}^{s}\mathbb{E}[|Y^{\eta}_{t}|^{2}]\mathrm{d}t\leq\mathbb{E}_{\mu}[\mathscr{H}(\pi^{\mu}(\cdot|X)|\pi^{\nu}(\cdot|X))]\leq\frac{\Lambda(\varphi^{\nu}_{0})}{2T}\,\mathbf{W}_{2}^{2}(\mu,\nu)\,.$
We conclude this section with one last technical bound.
Proposition 2.5.
@@
In view of Girsanov’s Theorem identity (2.4), it is enough to notice that
$2\mathscr{H}(\mathcal{L}(X^{\psi^{\mu},\rho}_{[0,\delta T]})|\mathcal{L}(X^{\psi^{\nu},\rho}_{[0,\delta T]}))\leq\int_{0}^{T}\mathbb{E}[|Y^{\eta}_{s}|^{2}]\mathrm{d}s+2\mathscr{H}(\mathcal{L}(X^{\varphi^{\mu},\mu}_{(1-\delta)T})|\mathcal{L}(X^{\varphi^{\nu},\nu}_{(1-\delta)T}))$
[…]
where we have relied on a second application of Girsanov’s Theorem (as we did for (2.3)), combined with the time-reversal identities (1.6). Applying Theorem 2.1 and Section 2.1 we concludes our proof. ∎
3. Proofs of the main results
@@
Assume H 1 and H 2, fix $\delta\in(0,1)$, and let $C^{\psi^{\nu}}_{\delta}\coloneqq T\biggl(\int_{0}^{\delta T}e^{\int_{0}^{s}2\lambda(\psi^{\nu}_{t})\mathrm{d}t}\mathrm{d}s\biggr)^{-1}$. Then we have
@@
As a corollary, if we define $C_{\rho\nu}^{\delta}\coloneqq C^{\psi^{\nu}}_{\delta}\,\biggl(\frac{\delta}{1-\delta}+3\,\Lambda(\varphi^{\nu}_{0})+2\sqrt{\Lambda(\varphi^{\nu}_{0})\,C^{\varphi^{\nu}}}\biggr)$, then we have
@@
From Itô’s formula and (2.1b), for all $s\leq\delta T$ we have
@@
which combined with Grönwall’s lemma gives $\mathbb{E}[|Y^{\theta}_{0}|^{2}]\,e^{\int_{0}^{s}2\lambda(\psi^{\nu}_{t})\mathrm{d}t}\leq\mathbb{E}[|Y^{\theta}_{s}|^{2}]$, that integrated over $s\in[0,\delta T]$ reads as
@@
Next, notice that from Girsanov’s theory (namely, the energy entropy identity (2.4)) we may recognize in the above right-hand side the relative entropy on the path space between the Schrödinger bridge from $\rho$ to $\mu$ and the Schrödinger bridge from $\rho$ to $\nu$, restricted on the time interval $0,\delta T]$, that is
@@
By recalling the time-reversal identities (1.6) and by applying the disintegration property of relative entropies (cf. [Nut21, Lemma 1.6] and [Léo14, Appendix A]) and Girsanov’s Theorem (w.r.t. the backward corrector process $Y^{\eta}_{\cdot}$) we deduce that
@@
This proves our first claim. By recalling the identity (2.2) we then have
@@
which can be bounded with Section 2.1 and with Theorem 2.1, yielding to our second claim. ∎
In theAppendix we specify Theorem 3.1 to diverse settings and in Corollaries B.1 and B.2 there we prove the asymptotic bounds stated in Theorem 1.1 in the Introduction.